# CooperTrim: Adaptive Data Selection for Uncertainty-Aware Cooperative Perception

### v2

@@ 1 Introduction
Cooperative perception (Wang et al., 2020; Xu et al., 2023a) enables autonomous agents to exchange information about invisible or uncertain regions to enhance a range of tasks, including detection (Qiu et al., 2022), prediction (Wang et al., 2025), mapping (Ahmad et al., 2020), and navigation (Cui et al., 2022). The benefits of additional vantage points come at the cost of communication. Balancing bandwidth overhead and accuracy, modern approaches adopt intermediate (over early and late) fusion (Wang et al., 2020; Xu et al., 2023a) to share task-oriented lightweight encoded representations of the environment. To further improve communication efficiency in intermediate fusion, three strategies have been proposed: compression (Wang et al., 2020; Xu et al., 2022a), which reduces the size of the features but risks information loss if targeted at higher compression ratio (i.e., lossy compression); selection (Hu et al., 2023; Liu et al., 2020), which transmits only useful data or selects agents; and hybrid (Yuan et al., 2022; Yang et al., 2023), combining both for optimal bandwidth reduction. For example, Where2comm (Hu et al., 2022) uses threshold-based spatial confidence maps for feature selection, UniSense (Ren et al., 2025) focuses on uncertainty-driven data exchange, and SwissCheese (Xie et al., 2024) exploits the disparity in semantic information on features between different spatial regions on different channels to perform fixed threshold-based selection. While these approaches reduce the exchange volume, the bandwidth demands still strain wireless technologies (Mo et al., 2025).
@@ 2 Related Work
Cooperative Perception. Resilient operation of autonomous agents depends on their onboard perception, which is often limited by blind spots and uncertainties. A range of cooperative perception mechanisms have been proposed, from early fusion and edge assistance (Zhang et al., 2021) sharing raw sensor data, to late fusion of regional processing results while missing holistic scene details. The most prominent are intermediate feature fusion methods (Chen et al., 2019), however, there remains a gap between bandwidth demand and availability. V2VNet (Wang et al., 2020) uses compressed LiDAR BEV features and GNN aggregation, yet often transmitting redundant data. AttFuse (Xu et al., 2022b) and DiscoNet (Li et al., 2021) offer attentive fusion and collaboration graphs, but lack intelligent data selection. CoBEVT (Xu et al., 2023a) integrates multi-vehicle BEV features via SinBEVT and FuseBEVT for segmentation, yet struggles with inefficient feature compression. Unlike these approaches, CooperTrim selectively transmits essential perception features, reducing bandwidth usage, improving network efficiency, and maintaining accuracy under constraints.
@@ 3.1 Problem Statement
We address the challenge of selective feature sharing for cooperative perception tasks. Rather than designing a new model, we aim at a selection framework that is applicable to a set of similar feature-sharing mechanisms (Xu et al. (2023a); Hu et al. (2022); Li et al. (2021). Given a set of features $\mathcal{F}=\{f_{1},f_{2},\dots,f_{n}\}$ extracted from input data, where $n$ is the number of feature channels in latent representations, our objective is to identify a subset $\mathcal{S}\subseteq\mathcal{F}$ such that the number of selected features $|\mathcal{S}|$ is minimized, while the accuracy of the perception task, denoted as a function $A(\mathcal{S})$ of the selected features, is maximized. Formally, the optimization problem can be expressed as $\min_{\mathcal{S}\subseteq\mathcal{F}}{|\mathcal{S}|},\max_{\mathcal{S}\subseteq\mathcal{F}}{A(\mathcal{S})}$. This bi-objective optimization problem poses significant challenges due to the inherent tension between feature abundance and bandwidth limitations.
@@ 3.3 Training
Training Strategy. For training the loss function in CooperTrim, we use an $\epsilon$-Greedy method inspired by the reinforcement learning $\epsilon$-greedy exploration strategy (Liu et al., 2022). However, we make some adaptations for our training purpose. With $\epsilon$ probability we request entire feature set and with $(1-\epsilon)$ probability we exploit the knowledge. By exploiting the knowledge, we request the data as per learned thresholds (i.e. conformal predictions’s quantile threshold $q$ and cross-attention module’s mask threshold $\tau$). We see considerable improvement in bandwidth requirement with this training method. Occasional updates considering all features can stabilize the optimization trajectory by smoothing out erratic updates caused by noise from partial features. This can prevent the model from diverging or oscillating around suboptimal points, leading to smoother convergence to a better solution. The $\epsilon$-Greedy method balances exploration (full data $D_{\text{full}}$ with probability $\epsilon$) and exploitation (partial data $D_{\text{partial}}$ with probability $(1-\epsilon)$). This balance stabilizes optimization by reducing gradient noise through periodic full-data updates. As formalized in Theorem 1, this strategy reduces both the bias and variance of the gradient estimator compared to using only partial data. The expected gradient
@@
ensures smoother convergence while adhering to bandwidth constraints. We formalize the effectiveness of the $\epsilon$-Greedy training strategy in the following theorem.
@@
Theorem 1 (Effectiveness of $\epsilon$-Greedy Training).
@@ 4 Experiments
We conduct six experiments to demonstrate CooperTrim’s efficiency in task performance and bandwidth usage. Referred to as CooperTrim, we apply CooperTrim to CoBEVT (Xu et al., 2023a) for evaluation, focusing on semantic segmentation on the OPV2V (Xu et al., 2022b) dataset and 3D detection on OPV2V and V2V4Real (Xu et al., 2023b) datasets.. We choose CoBEVT for its robust performance in cooperative perception compared to existing works. CooperTrim is the first to address selective cooperative perception in semantic segmentation.
@@
Figure 3: Across-the-board Bandwidth Comparison. C: Compression. FS: Feature Selection. AS: Agent Selection. CooperTrim consumes the lowest bandwidth among baselines.
@@
Post-selection Compression Compatibility.
Dynamic
Lane
Road
Compression-based methods like CoBEVT (Xu et al., 2023a) and AttFuse (Xu et al., 2022b) reduce data size for network efficiency. Yet, CooperTrim’s data selection enhances efficiency by compressing post-selection, optimizing bandwidth. We apply lossy quantization to reduce precision and volume, then lossless methods for further reduction. Figure 4 compares IoU and bandwidth for CooperTrim, CoBEVT (Xu et al., 2023a), and AttFuse (Xu et al., 2022b) at 1x, 8x, and 32x rates for the segmentation task. CooperTrim excels in IoU and bandwidth savings, notably at 32x (1.46% vs. 3.76% for CoBEVT and 10.62% for AttFuse in 32x Dyn) [See Appendix A.6 for detailed results].
@@
Figure 5: Increased data requests align with higher scene complexity. For dynamic objects, complexity in number and positioning rises in Frames 1200, 200, and 1700. For static elements, complexity grows in Frames 900, 250, and 1600, with more intersections and lane orientations. Visualizations on the right highlight the frames at vertical lines; Frame 1200, 200, 1700 (dynamic) and 900, 250, 1600 (static). Green dashed lines show baseline CoBEVT IoU. Green solid lines show CooperTrim IoU.
Environment Adaptation. Figure 5 shows the inference results for CooperTrim over multiple frames. We see variations in bandwidth implying variable feature request in each frame. The figure shows increased data requests correlate with higher scene complexity. In segmenting the dynamic objects, an increase in complexity can be referred to criticality in positioning of an increased amount of vehicles and traffic participants. Similarly, an increase in complexity in the static task can be referred to an increase in the number of intersections or lane orientations in the scene. Figure 5 shows that in both cases, i.e., increased amount of vehicles, and increased topological complexity in the roadways, CooperTrim accordingly shares more data to preserve task performance, whereas in other cases, gracefully adapts to request less sharing, saving the precious bandwidth to other agents in need. In critical frame ranges such as 1000-1800 (Dynamic), 1000-1500 (Lane), and 1200-1400 (Road), where baseline CoBEVT (Xu et al., 2023a) consistently underperforms, CooperTrim achieves much higher IoU, illustrating the effectiveness of its adaptive threshold masking mechanism in prioritizing key features and managing scene complexity.
@@ A.9 Conformal Temporal Uncertainty
Figure 13: Conformal Temporal Uncertainty

### v3

# arXiv:2602.13287v3[cs.CV] 27 Feb 2026

@@ 1 Introduction
Cooperative perception (Wang et al., 2020; Xu et al., 2023a) enables autonomous agents to exchange information about invisible or uncertain regions to enhance a range of tasks, including detection (Qiu et al., 2022), prediction (Wang et al., 2025), mapping (Ahmad et al., 2020), and navigation (Cui et al., 2022; 2026). The benefits of additional vantage points come at the cost of communication. Balancing bandwidth overhead and accuracy, modern approaches adopt intermediate (over early and late) fusion (Wang et al., 2020; Xu et al., 2023a) to share task-oriented lightweight encoded representations of the environment. To further improve communication efficiency in intermediate fusion, three strategies have been proposed: compression (Wang et al., 2020; Xu et al., 2022a), which reduces the size of the features but risks information loss if targeted at higher compression ratio (i.e., lossy compression); selection (Hu et al., 2023; Liu et al., 2020), which transmits only useful data or selects agents; and hybrid (Yuan et al., 2022; Yang et al., 2023), combining both for optimal bandwidth reduction. For example, Where2comm (Hu et al., 2022) uses threshold-based spatial confidence maps for feature selection, UniSense (Ren et al., 2025) focuses on uncertainty-driven data exchange, and SwissCheese (Xie et al., 2024) exploits the disparity in semantic information on features between different spatial regions on different channels to perform fixed threshold-based selection. While these approaches reduce the exchange volume, the demand still strains wireless bandwidth (Mo et al., 2025).
@@ 2 Related Work
Cooperative Perception. Resilient operation of autonomous agents depends on their onboard perception, which is often limited by blind spots and uncertainties. A range of cooperative perception mechanisms have been proposed, from early fusion and edge assistance (Zhang et al., 2021; Zhu et al., 2024) sharing raw sensor data, to late fusion of regional processing results while missing holistic scene details. The most prominent are intermediate feature fusion methods (Chen et al., 2019), however, there remains a gap between bandwidth demand and availability. V2VNet (Wang et al., 2020) uses compressed LiDAR BEV features and GNN aggregation, yet often transmitting redundant data. AttFuse (Xu et al., 2022b) and DiscoNet (Li et al., 2021) offer attentive fusion and collaboration graphs, but lack intelligent data selection. CoBEVT (Xu et al., 2023a) integrates multi-vehicle BEV features via SinBEVT and FuseBEVT for segmentation, yet struggles with inefficient feature compression. Unlike these approaches, CooperTrim selectively transmits essential perception features, reducing bandwidth usage, improving network efficiency, and maintaining accuracy under constraints.
@@ 3.1 Problem Statement
We address the challenge of selective feature sharing for cooperative perception tasks. Rather than designing a new model, we aim at a selection framework that is applicable to a set of similar feature-sharing mechanisms (Xu et al. (2023a); Hu et al. (2022); Li et al. (2021). Given a set of features $\mathcal{F}=\{f_{1},f_{2},\dots,f_{n}\}$ extracted from input data, where $n$ is the number of feature channels in latent representations, our objective is to identify a subset $\mathcal{S}\subseteq\mathcal{F}$ such that the number of selected features $|\mathcal{S}|$ is minimized, while the accuracy of the perception task, denoted as a function $A(\mathcal{S})$ of the selected features, is maximized. Formally, the optimization problem can be expressed as $\min_{\mathcal{S}\subseteq\mathcal{F}}{|\mathcal{S}|},\max_{\mathcal{S}\subseteq\mathcal{F}}{A(\mathcal{S})}$. This bi-objective optimization problem is challenging due to the inherent tension between feature abundance and bandwidth limitations.
@@ 3.3 Training
Training Strategy. For training the loss function in CooperTrim, we use an $\epsilon$-Greedy method inspired by the reinforcement learning $\epsilon$-greedy exploration strategy (Liu et al., 2022). However, we make some adaptations for our training purpose. With $\epsilon$ probability we request entire feature set and with $(1-\epsilon)$ probability we exploit the knowledge. By exploiting the knowledge, we request the data as per learned thresholds (i.e. conformal predictions’s quantile threshold $q$ and cross-attention module’s mask threshold $\tau$). We see considerable improvement in bandwidth requirement with this training method. Occasional updates considering all features can stabilize the optimization trajectory by smoothing out erratic updates caused by noise from partial features. This can prevent the model from diverging or oscillating around suboptimal points, leading to smoother convergence to a better solution. The $\epsilon$-Greedy method balances exploration (full data $D_{\text{full}}$ with probability $\epsilon$) and exploitation (partial data $D_{\text{partial}}$ with probability $(1-\epsilon)$). This balance stabilizes optimization by reducing gradient noise through periodic full-data updates. As formalized in Proposition 1, this strategy reduces both the bias and variance of the gradient estimator compared to using only partial data. The expected gradient
@@
ensures smoother convergence while adhering to bandwidth constraints. We formalize the effectiveness of the $\epsilon$-Greedy training strategy in the following proposition.
@@
Proposition 1 (Effectiveness of $\epsilon$-Greedy Training).
@@ 4 Experiments
We conduct six experiments to demonstrate CooperTrim’s efficiency in task performance and bandwidth usage. Referred to as CooperTrim, we apply CooperTrim to CoBEVT (Xu et al., 2023a) for evaluation, focusing on semantic segmentation on the OPV2V (Xu et al., 2022b) dataset and 3D detection on OPV2V and V2V4Real (Xu et al., 2023b) datasets. We choose CoBEVT for its robust performance in cooperative perception compared to existing works. CooperTrim is the first to address selective cooperative perception in semantic segmentation.
@@
Post-selection Compression Compatibility. Compression-based methods like CoBEVT (Xu et al., 2023a) and AttFuse (Xu et al., 2022b) reduce data size for network efficiency. Yet, CooperTrim’s data selection enhances efficiency by compressing post-selection, optimizing bandwidth. We apply lossy quantization to reduce precision and volume, then lossless methods for further reduction. Figure 3 compares IoU and bandwidth for CooperTrim, CoBEVT (Xu et al., 2023a), and AttFuse (Xu et al., 2022b) at 1x, 8x, and 32x rates for the segmentation task. CooperTrim excels in IoU and bandwidth savings, notably at 32x (1.46% vs. 3.76% for CoBEVT and 10.62% for AttFuse in 32x Dyn) [See Appendix A.6 for detailed results].
Dynamic
Lane
Road
Environment Adaptation. Figure 4 shows the inference results for CooperTrim over multiple frames. We see variations in bandwidth implying variable feature request in each frame. The figure shows increased data requests correlate with higher scene complexity. In segmenting the dynamic objects, an increase in complexity can be referred to criticality in positioning of an increased amount of vehicles and traffic participants. Similarly, an increase in complexity in the static task can be referred to an increase in the number of intersections or lane orientations in the scene. Figure 4 shows that in both cases, i.e., increased amount of vehicles, and increased topological complexity in the roadways, CooperTrim accordingly shares more data to preserve task performance, whereas in other cases, gracefully adapts to request less sharing, saving the precious bandwidth to other agents in need. In critical frame ranges such as 1000-1800 (Dynamic), 1000-1500 (Lane), and 1200-1400 (Road), where baseline CoBEVT (Xu et al., 2023a) consistently underperforms, CooperTrim achieves much higher IoU, illustrating the effectiveness of its adaptive threshold masking mechanism in prioritizing key features and managing scene complexity (see Appendix A.5).
@@
Figure 4: Increased data requests align with higher scene complexity. For dynamic objects, complexity in number and positioning rises in Frames 1200, 200, and 1700. For static elements, complexity grows in Frames 900, 250, and 1600, with more intersections and lane orientations. The right part highlights the frames at vertical lines; Frame 1200, 200, 1700 (dynamic) and 900, 250, 1600 (static). Green dashed lines: baseline CoBEVT IoU. Green solid lines: CooperTrim IoU.
@@
Figure 8: Across-the-board Bandwidth Comparison. C: Compression. FS: Feature Selection. AS: Agent Selection. CooperTrim consumes the lowest bandwidth.
@@ 5 Conclusion
Acknowledgement
This research was sponsored by the OUSD (R&E)/RT&L and was accomplished under Cooperative Agreement Number W911NF-20-2-0267. The views and conclusions contained in this document are those of the authors and should not be interpreted as representing the official policies, either expressed or implied, of the ONR, ARL and OUSD(R&E)/RT&L or the U.S. Government. The U.S. Government is authorized to reproduce and distribute reprints for Government purposes notwithstanding any copyright notation herein.
@@ A.9 Conformal Temporal Uncertainty
Figure 13: CooperTrim estimates conformal temporal uncertainty by comparing the current frame’s encoded features with the ego vehicle’s fused features from recent frames, highlighting what has changed over time (Non-Conformity Scores). It then uses a learnable quantile-based gating rule to identify high-deviation (more uncertain) features as the most relevant candidates for request.